jhonyy9 (jhonyy9):

how can proving this ? e^(pi*i) = -1

jhonyy9 (jhonyy9):

so this mean that e^(pi*i) = i^2 make ln of both sides ln e^(pi*i) = ln i^2 (pi*i)ln e = 2ln i pi*i = 2ln i divide both sides by 2 (pi*i)/2 = ln i can being this really true that logarithm natural of ,,i" is equal (i*pi)/2 ?

jhonyy9 (jhonyy9):

@TheSmartOne @Kainui @mathmate @ganeshie8 @welshfella

OpenStudy (mathmate):

Use the expansion of trigonometric and exponential functions. \(e^{x}=1+x/1!+x^2/2!+...\) \(e^{i\pi}=1+(i\pi)/1!+(i\pi)^2/2!+...\) \(=1+i^2(\pi)^2/2!+...+(i\pi)/1!+(i\pi)^3/3!+...\) \(=1-(\pi)^2/2!+...+i(\pi)/1!-i(\pi)^3/3!+...\) \(=cos(\pi)+(i)sin(\pi)\) \(=-1+0\) \(=-1\)

jhonyy9 (jhonyy9):

ty. @mathmate this is really GREAT - thank you very much !!!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (zzr0ck3r):

Infinite sums is the only way I have seen it derived.

jhonyy9 (jhonyy9):

hi

Parth (parthkohli):

hello

jhonyy9 (jhonyy9):

so you accept it now how is here proven ?

Parth (parthkohli):

It is correct that \(e^{i\pi} = -1\), but it is not true that \(\ln(-1) = i\pi\), because it is not unique. I could also have said that \(\ln(-1) = 3i\pi = 5i\pi = \cdots\) So because of domain problems, we only think about logarithms in real numbers. Otherwise, it leads to very contradictory results.

jhonyy9 (jhonyy9):

but in place of (-1) i used i^2

Parth (parthkohli):

So exponential functions in complex numbers are OK because they are unique, but logarithmic functions cannot be used at the elementary level.

Parth (parthkohli):

Here is another way to think why logarithms should not be used for complex numbers: We know that \(i = 1i = i^4i = i^5\). Now because \(i = i^5\) we can say that \(\ln(i) = \ln(i^5)\). So \(\ln(i) = 5 \ln(i)\) and finally \(\ln(i) = 0\). But \(e^0 \ne i\) so this is false.

jhonyy9 (jhonyy9):

can i reposting these all on OS newly please - bc. this your opinion about this is very very interesting

jhonyy9 (jhonyy9):

Parth this is possible or not reposting it ?

Parth (parthkohli):

Yes, please do.

jhonyy9 (jhonyy9):

ok. but how ?

Parth (parthkohli):

I can do it for you.

jhonyy9 (jhonyy9):

can you tel me how ?bc. this is ohhh. ok please do it than now

jhonyy9 (jhonyy9):

thank you very much